OpenStudy (anonymous):

i need help!!! the table below represents the distance of a car from its destination as a function of time time (hours) distance (miles) 0 900 1 850 2 800 3 750 what is the y intercept of a function and what does this tell you about the car? calculate the average rate of change of the function represented by the table between x=1 to x=3? what would be the domain of this function if the car traveled the same rate until it reached its destination?

Vocaloid (vocaloid):

start by finding the y-intercept

Vocaloid (vocaloid):

if x = time and y = distance, what is the distance when x = 0 ?

OpenStudy (anonymous):

900

OpenStudy (anonymous):

sorry i had to take my dog out

Vocaloid (vocaloid):

good, so the y-intercept is (0,900) which means the car starts off at 900 miles away from its destination

Vocaloid (vocaloid):

for average rate of change, you want to find (distance when x = 3) - (distance when x = 1)

Vocaloid (vocaloid):

then divide that by 2

OpenStudy (anonymous):

distance equals 750 when time is 3 and the distance equals 850 when the time equals one

Vocaloid (vocaloid):

good, then we subtract

Vocaloid (vocaloid):

750 - 850 = -100

Vocaloid (vocaloid):

then divide by 2, to get -50

OpenStudy (anonymous):

okay

Vocaloid (vocaloid):

the slope is -50 and the y-intercept is 900 so we have the equation y = -50x + 900 to find the domain, let y = 0 and solve for x 0 = -50x + 900 x = ?

OpenStudy (anonymous):

so 0-900?

Vocaloid (vocaloid):

-900 = -50x x = ?

OpenStudy (anonymous):

okay so -900/-50

Vocaloid (vocaloid):

right, keep going

Vocaloid (vocaloid):

-900/-50 = ??

OpenStudy (anonymous):

18

OpenStudy (anonymous):

right?

Vocaloid (vocaloid):

right, so the domain is all real numbers from x = 0 to x = 18 you can write this as [0,18]

OpenStudy (anonymous):

okay

Vocaloid (vocaloid):

and that's it

OpenStudy (anonymous):

thank you so much